Nursery manufactures and sells pots. Currently, 250 000 units are sold per year at $3.20 per unit. Fixed costs are $190 000 per year. Variable costs are $1.30 per unit.

Required
a. What is the current annual profit? (1 mark)

b. What is the present break-even point in revenues? (3 marks)

Calculate the new profit for each of the following changes:
c. a $0.30 per unit increase in variable costs (1 marks)
d. a 7% increase in fixed costs and a 7% increase in units sold (3 marks)
e. a 12% decrease in fixed costs, a 12% decrease in selling price, a 12% decrease in variable cost per unit and a 30% increase in units sold (4 marks)

Calculate the new break-even point in units for each of the following changes:f. a 10% increase in fixed costs (1 marks)
g. a 10% increase in selling price and a $20 000 increase in fixed costs (2 marks)
Answer #1

a)

Current annual profit = Units sold * (Unit price - Variable cost) - Fixed cost

= 250000*(3.2-1.3)-190000

= $ 285,000

b)

Contribution margin ratio = (3.2-1.3)/3.2 = 0.59375

Present break-even point in revenues = Fixed cost / Contribution margin ratio

= 190000 / 0.59375

= $ 320,000

c)

New variable csot = $ 1.3+0.3 = $ 1.6

New annual profit = Units sold * (Unit price - Variable cost) - Fixed cost

= 250000*(3.2-1.6)-190000

= $ 210,000

d)

New fixed cost = 190000*1.07 = $ 203,300

Units sold = 250000*1.07 = 267,500 units

New annual profit = 267500*(3.2-1.3)-203300

= $ 304,950

e)

New fixed cost = 190000*(1-0.12) = $ 167,200

New selling price = 3.2*(1-0.12) = $ 2.816

New variable cost = 1.3*(1-0.12) = $ 1.144

Units sold = 250000*(1+0.3) = 325,000 units

New annual profit = 325000*(2.816-1.144)-167200

= $ 376,200

f)

New fixed cost = 190000*(1+0.1) = $ 209,000

New break-even point in units = New fixed cost / (Selling price - Variable cost)

= 209000/(3.2-1.3)

= 110,000 units

g)

New selling price = 3.2*(1+0.1) = $ 3.52

New fixed costs = 190000+20000 = $ 210,000

New break-even point in units = New fixed cost / (Selling price - Variable cost)

= 210000/(3.52-1.3)

= 94,595 units